# Measuring the Impact of Missingness in Traffic Stop Data

### v1

@@ Appendix A SOPP Data Summary
CO Aurora 2012 - 2020 257606 16 0.146 NA NA
CO Denver 2010 - 2018 1870731 14 NA NA NA
CO Statewide 2010 - 2017 3112853 20 12.903 7.424 0.000
@@
OH Cincinnati 2009 - 2018 315281 25 0.000 NA NA
OH Columbus 2012 - 2016 128157 18 0.000 0.000 NA
OH Statewide 2010 - 2017 7225577 24 8.704 0.000 80.163
@@
WA Seattle 2006 - 2015 319959 27 99.896 NA NA
WA Statewide 2009 - 2018 11333425 30 28.176 0.000 0.000
WA Tacoma 2007 - 2017 271912 15 NA NA NA
WI Madison 2007 - 2020 331450 25 1.122 NA NA
WI Statewide 2010 - 2016 1058902 44 14.505 0.000 0.000
@@ Appendix B Causal DAG
We introduce three indicator variables $R_{X_{2}}$, $R_{M}$, and $R_{Y}$ to denote whether an instance contains missing data for partially-observed covariates ($X_{2}$), the treatment ($M$), or outcome ($Y$). Moreno-Betancur et al., (2018) also include unmeasured causes of missingness ($W$) that might affect the probability of missingness. $W$ could include factors such as the weather or departmental policy changes for the precinct that might affect data records.
Having no arrows to the missingness indicators, $M$, would mean the missingness mechanism is MCAR and entirely dictated by unmeasured factors, $W$. However, in the case of SOPP data, we notice that complete and missing covariates have the potential to affect the missingness of the stop, the outcome, and the covariates themselves as discussed in Section 3. For example, a certain police department might have higher rates of missing data in their reporting (Pierson et al.,, 2020) or officers might omit data at differing rates for different races (Chanin and Welsh,, 2021). Hence, Figure 3 contains arrows from $X_{1}$ and $X_{2}$ to the three missingness indicators.
Additionally, policing datasets such as SOPP inherently only include observations with one level of the treatment - individuals who were stopped ($M=1$) - and contain no report on the number of individuals who were observed but not stopped ($M=0$) (Knox et al.,, 2020).
@@
$\mathbf{X_{2}}$ More stops among Black civilians More searches among Black civilians More race missingness for Black stops (Chanin and Welsh,, 2021) Less stops among white civilians Complacency in search records for some race
@@ D.2 Extreme Cases
Figure 19: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for CO Statewide, the disparity estimates can change direction from discrimination against Black drivers to the opposite direction.
Figure 20: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for MD Statewide, the ATE bounds are always above zero, even in the most extreme cases, indicating racial discrimination against Black drivers.
Figure 21: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for WI Statewide, the ATE bounds are almost always above zero, indicating discrimination against Black drivers.

### v2

@@ Appendix A SOPP Data Summary
CO, Aurora 2012 - 2020 257606 16 0.146 NA NA
CO, Denver 2010 - 2018 1870731 14 NA NA NA
CO, Statewide 2010 - 2017 3112853 20 12.903 7.424 0.000
@@
OH, Cincinnati 2009 - 2018 315281 25 0.000 NA NA
OH, Columbus 2012 - 2016 128157 18 0.000 0.000 NA
OH, Statewide 2010 - 2017 7225577 24 8.704 0.000 80.163
@@
WA, Seattle 2006 - 2015 319959 27 99.896 NA NA
WA, Statewide 2009 - 2018 11333425 30 28.176 0.000 0.000
WA, Tacoma 2007 - 2017 271912 15 NA NA NA
WI, Madison 2007 - 2020 331450 25 1.122 NA NA
WI, Statewide 2010 - 2016 1058902 44 14.505 0.000 0.000
@@ Appendix B Causal DAG
We introduce three indicator variables $R_{X_{2}}$, $R_{M}$, and $R_{Y}$ to denote whether an instance contains missing data for partially-observed covariates ($X_{2}$), the treatment ($M$), or outcome ($Y$). moreno2018canonical also include unmeasured causes of missingness ($W$) that might affect the probability of missingness. $W$ could include factors such as the weather or departmental policy changes for the precinct that might affect data records.
Having no arrows to the missingness indicators, $M$, would mean the missingness mechanism is MCAR and entirely dictated by unmeasured factors, $W$. However, in the case of SOPP data, we notice that complete and missing covariates have the potential to affect the missingness of the stop, the outcome, and the covariates themselves as discussed in Section 3. For example, a certain police department might have higher rates of missing data in their reporting (pierson2020large) or officers might omit data at differing rates for different races (chanin2021). Hence, Figure 3 contains arrows from $X_{1}$ and $X_{2}$ to the three missingness indicators.
Additionally, policing datasets such as SOPP inherently only include observations with one level of the treatment - individuals who were stopped ($M=1$) - and contain no report on the number of individuals who were observed but not stopped ($M=0$) (knox2020administrative).
@@
$\mathbf{X_{2}}$ More stops among Black civilians More searches among Black civilians More race missingness for Black stops (chanin2021) Less stops among white civilians Complacency in search records for some race
@@ D.2 Extreme Cases
Figure 19: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for CO, Statewide, the disparity estimates can change direction from discrimination against Black drivers to the opposite direction.
Figure 20: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for MD, Statewide, the ATE bounds are always above zero, even in the most extreme cases, indicating racial discrimination against Black drivers.
Figure 21: Dashed line at zero indicates no differential racial impact. Disparity is the difference in proportion of search for Black minus proportion of search for white. Disparity is measured as bounds on the search ATE, given recorded outcome, for different levels of the $\rho$ parameter (the overall proportion of racially discriminatory stops). The $x$-axis of each plot presents the ignore NA case as well as the two most extreme scenarios of allocating NA race observations to either Black or white. Note that for WI, Statewide, the ATE bounds are almost always above zero, indicating discrimination against Black drivers.